OpenStudy (catpidigm):

Fan and Medal! Which linear function includes the points (-3, 1) and (-2, 4)? A. y = 3x +10 B. y = 3x-6 C. y = 1/3x + 2 D. y = -3x + 1 I'm taking a test, so I will probably be posting alot of questions.

OpenStudy (will.h):

The slope would determine that Find the slope Slope = Y2-Y1/X2-X1 Slope = 4-1/-2+3 Slope = 3/1 Slope = 3 Therefore the answer with slope 3 is A

OpenStudy (catpidigm):

Thank you!

OpenStudy (will.h):

You welcome